OpenStudy (usukidoll):

My first probability question :P In how many ways can you choose a group of three people from a group of six?

OpenStudy (usukidoll):

(will type attempt in a bit)

OpenStudy (usukidoll):

\[\frac{n!}{r!(n-r)!}\] so... seeing that this is a combination type problem, I have to use that formula if I let n = 6 and r = 3, then \[\frac{6!}{3!(6-3)!}\] \[\frac{6!}{3!(3)!}\] \[\frac{720}{36} = 20\] so there's 20 different ways to choose a group of 3 from 6.

OpenStudy (usukidoll):

I think that's right???? I mean the combination formula is straight forward.

OpenStudy (usukidoll):

@agent0smith geez man...how long has it been since I saw you on OpenStudy?

OpenStudy (jameshorton):

can somebody help me on mines

OpenStudy (usukidoll):

oh yeah like posting your question on mine will help smh.

OpenStudy (freckles):

that seems right "In how many ways can you choose a group of three people from a group of six? " There is 6 ways to choose the first There is 5 ways to choose the second There are 4 ways to choose the third and then we need to figure out how we can arrange the 3 objects so the above product over 3! \[\frac{6 \cdot 5 \cdot 4}{3!} \text{ which is the same as } \frac{6 \cdot 5 \cdot 4 \cdot 3!}{3! \cdot 3!} =\frac{6!}{3! \cdot 3!}\]

OpenStudy (usukidoll):

ah thanks. I haven't touch probability in forever which is why I'm shaky at it.

OpenStudy (freckles):

I can understand why you would be shaky. It is hard to count especially when the numbers get really big.

OpenStudy (freckles):

And the orders get crazy.

OpenStudy (usukidoll):

yeah... but I haven't even touch this subject at all. xD! Apparently Discrete Math has this... >_<

OpenStudy (usukidoll):

so it will be like 6! = 6 x 5 x 4 x 3 x 2 x 1 over 3! = 3 x 2 x 1 and at least the 3 x 2 x 1 cancels 6 x 5 x 4 / 3 x 2 x 1 = 120/6=20.. ah I see ...